OpenStudy (haley.fieldss):

Can some one simply 6r+r-5r

OpenStudy (maddielb):

2r

KjSaif (kjsaif):

explain @MaddieLB

likeabossssssss (likeabossssssss):

^^

OpenStudy (maddielb):

You add r to 6r which gives you 7r then subtract 5r from 7r and get 2r

OpenStudy (haley.fieldss):

thanks b

KjSaif (kjsaif):

6r+r−5r =6r+r+−5r Combine Like Terms: =6r+r+−5r =(6r+r+−5r) =2r

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (haley.fieldss):

can some one simply 5x+ 2 (x+6)?

likeabossssssss (likeabossssssss):

plz put it in a new post thank u

OpenStudy (maddielb):

^

KjSaif (kjsaif):

^^^^^^^

likeabossssssss (likeabossssssss):

4 this q dont come spam

KjSaif (kjsaif):

^^

likeabossssssss (likeabossssssss):

:)